OpenStudy (richgall____):

yall are no help i just need help on 1 problem and no one knows it :/

OpenStudy (anonymous):

What's the problem?

OpenStudy (richgall____):

Gina looks at the architectural plan of a four-walled room in which the walls meet each other at right angles. The length of one wall in the plan is 17 inches. The length of the diagonal of the floor of the room in the plan is approximately 18.79 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work.

OpenStudy (anonymous):

I find drawing out problems like this tends to help 

OpenStudy (anonymous):

I would think that you would use the Patagonian theorem to solve for the missing side, and that might answer your question. If a^2+b^2=c^2 and c is the diagonal, then c^2-a^2=b^2

OpenStudy (anonymous):

^ The above is correct. To prove this is a square, you need to prove that the perpendicular side is 17 inches. You have two lengths of a right angled triangle, so Pythagoras' theorem is the way to go. c^2 (the diagonal squared) = a^2 (one length squared) + b^2 (the other length squared) Try substituting your values into that equation

OpenStudy (richgall____):

thanks u guys